# On the Need for a Statistical Foundation in Scenario-Based Testing of Autonomous Vehicles

### v1

@@ IV-C Modelling Fidelity of Synthetic Data
Eq. (14) is a classical confidence statement [^7] This is not a Bayesian probability distribution of the actual error, but rather a frequentist confidence bound: if you repeated the process of testing and deriving the estimator, many times, the estimator would fall within the error range ($-\epsilon,+\epsilon$) around the true value with at least $1-\alpha$ probability. : with confidence at least $1-\alpha$, the synthetic-data estimate deviates [^8] In real safety cases, relative error often matters more than absolute error. from the real-data estimate by no more than $\epsilon$. Indeed, even if the estimates using real and synthetic data are close enough, they may still detect very different failures. Thus, our goal is not to replace existing fidelity definitions and studies like [23], but rather to complement them from an additional perspective.
@@
$|\hat{\theta}^{s}-\hat{\theta}^{r}|$ $=|\underbrace{(\hat{\theta}^{s}-\theta^{s})}_{\text{syn. samplingvar.}}\!\!+\underbrace{(\theta^{s}-\theta^{r})}_{\text{syn. bias }\Delta}+\!\!\!\underbrace{(\theta^{r}-\hat{\theta}^{r})}_{\text{real-w. sampling var.}}|$ (15)
The first and third terms representing sampling uncertainty can be reduced given more real and synthetic data collected. The second term $\Delta$ captures the simulator’s inherent quality, which is irreducible unless the simulator is reconfigured to be a better one. By understanding those three uncertainties, we propose a REF use case with the following steps:
@@
    Collect real-world data Execute a fixed number of real-world scenario tests within a given testing budget. Estimate $\hat{\theta}^{r}$ using a chosen estimator.
@@
    Generate synthetic data Use the current simulator configuration to run a number of synthetic scenario tests. Estimate $\hat{\theta}^{s}$ using the same estimator.
@@
    If REF certification fails, increase synthetic data Increase the number of simulator tests to reduce the variance in $\hat{\theta}^{s}$. Recompute REF. Repeat until: REF is satisfied, or additional samples yield no significant improvement.
@@
    If REF certification still fails, reconfigure simulator Modify the simulator to reduce the gap between real and synthetic performance (i.e., reduce bias $\Delta=\theta^{s}-\theta^{r}$). Repeat from Step 2.
@@
    If REF remains unsatisfied, quantify fidelity limit Given the current data, determine the smallest $\epsilon$ (and/or largest $\alpha$) for which REF is certified. The new $(\epsilon,\alpha)$ reflects the best fidelity we can certify with the available data.
@@
    Perform extended testing with certified simulator If REF is satisfied, scale up simulator testing to estimate $\theta^{r}$. Use certified $(\epsilon,\alpha)$ to bound uncertainty in final safety claims.
@@
    Step 4: Doubling $t_{s}$ to 4000 with $k_{s}=102$ yields $\Pr(|\Delta\hat{|}\leq 0.02)\approx 0.91<0.95$; $(0.02,0.05)$-REF still uncertified.
@@
    $\hat{\Delta}=-0.005,\,\,\sigma_{\hat{\Delta}}\approx 0.0089,\,\,\Pr(|\Delta\hat{|}\leq 0.02)>0.95,$
    so $(0.02,0.05)$-REF is certified with $\hat{\Delta_{c}}=-0.002$.
@@
    Applying the REF bias bound $\epsilon=0.02$, the adjusted 90% CI for $\theta^{r}$ becomes [^10] Proof sketch: We aim to construct a CI for $\theta^{r}$, accounting for the following two sources of uncertainty. Event A (with probability $\alpha_{A}$): the simulation CI misses $\theta^{s}$, $\theta^{s}\notin[I_{l},I_{u}]$, i.e., the $0.951-\alpha_{A}=$ 0.95 CI for $\hat{\theta}^{s}$ does not contain the true $\theta^{s}$; Event B (with probability $\alpha_{B}$): REF fails to bound the bias, $|\theta^{s}-\theta^{r}|>\epsilon$, i.e., the REF certification fails to ensure $|\theta^{s}-\theta^{r}|\leq\epsilon$. By the union bound, $P(A\cup B)\leq\alpha_{A}+\alpha_{B}$, it is evident that the probability that either the simulation CI misses $\theta^{s}$ or the REF certification fails is at most $\alpha_{a}+\alpha_{B}$. Thus, the probability that both the simulation CI and the REF certification are correct is at least $1-(\alpha_{A}+\alpha_{B})$. So, if neither A nor B fails (the simulation CI contains $\theta^{s}$ and the REF ensures $|\theta^{s}-\theta^{r}|\leq\epsilon$), then the adjusted CI (simulation CI expanded by $[I_{l}\!-\!\epsilon,I_{u}\!+\!\epsilon]$) contains $\theta_{r}$.:
    [0.026850.02, 0.02975+0.02]=[0.00685, 0.04975].0.026850.020.029750.020.006850.04975[0.02685-0.02,\ 0.02975+0.02]=[0.00685,\ 0.04975].[ 0.02685 - 0.02 , 0.02975 + 0.02 ] = [ 0.00685 , 0.04975 ] .
While the adjusted CI for $\theta_{r}$ may be (one-side) tighter than the CI directly calculated from real-world data along, it is certainly more actionable by certifying a smaller $\epsilon$ and/or adding more synthetic data. Enhanced statistical models are needed for Step 7 to capture more complexities, which is our future work. Furthermore, there are limitations like the failures are so rare that the sample means can no longer be approximated by normal distributions. We omit such cases, as our intent is to demonstrate the feasibility of REF, rather than provide a comprehensive formal treatment.
Another future direction is to define a Bayesian version of REF, where beliefs about the actual $\theta^{r}$ and $\theta^{s}$ are modeled, and being able to reason under rare failures. Once such a Bayesian REF is certified (e.g., with 95% probability that $\theta^{r}$ is no worse than $\theta^{s}$), existing models for “No Worse Than” reasoning [36, 43, 42, 46, 45] can be applied to support safety claims based on both real and simulated evidence.
@@ V Discussion and Conclusion
While we introduce pfs in our proof-of-concept models for scenario testing evidence, is pfs truly the best metric? Or should we consider higher abstraction level of “per trip” or “per mission”? What constitutes an acceptable pfs in AVs? Can we derive such thresholds from human-driven incident statistics at “per scenario” level? Answer to those question depends on how data are collected in practice for the specific AV under study, considering, e.g., if scenarios are i.i.d. and if pfs is constant for the data collected.
@@
Our goal in building a statistical foundation for scenario-based testing is not to replace existing approaches to AV safety, but to offer a complementary perspective: one that is explicitly driven by reasoning about residual operational risk. By mapping new challenges in AV testing to known problems in software reliability, we aim to bridge methodological gaps and enable more rigorous safety claims for AVs.

### v2

@@ IV-C Modelling Fidelity of Synthetic Data
Eq. (14) is a classical confidence statement [^8] This is not a Bayesian probability distribution of the actual error, but rather a frequentist confidence bound: if you repeated the process of testing and deriving the estimator, many times, the estimator would fall within the error range ($-\epsilon,+\epsilon$) around the true value with at least $1-\alpha$ probability. : with confidence at least $1-\alpha$, the synthetic-data estimate deviates [^9] In real safety cases, relative error often matters more than absolute error. from the real-data estimate by no more than $\epsilon$. Noted that, even if the estimates using real and synthetic data are close, they may still detect very different failures. Thus, our goal is not to replace existing fidelity definitions and studies like [23].
@@
$|\hat{\theta}^{s}-\hat{\theta}^{r}|$ $=|\underbrace{(\hat{\theta}^{s}-\theta^{s})}_{\text{sim. samplingvariance}}\!\!+\underbrace{(\theta^{s}-\theta^{r})}_{\text{sim. bias }\Delta}+\!\!\!\underbrace{(\theta^{r}-\hat{\theta}^{r})}_{\text{real-w. sampling variance}}|$
The first and third terms representing sampling uncertainty can be reduced given more real and synthetic data collected. The second term $\Delta$ captures the simulator’s inherent quality, which is irreducible unless the simulator is reconfigured to be a better one. By understanding those three uncertainties, we propose a REF use case with steps of:
@@
    Collect real-world data Test a fixed and limited number of scenarios using real-world data. Estimate $\hat{\theta}^{r}$.
@@
    Generate synthetic data Use the current simulator configuration to run a number of scenario tests. Estimate $\hat{\theta}^{s}$.
@@
    If REF certification fails, increase synthetic data to reduce the variance in $\hat{\theta}^{s}$. Recompute REF. Repeat until: REF is satisfied, or additional samples yield no improvement.
@@
    If REF certification still fails, reconfigure simulator to improve its accuracy, and hopefully reduce the gap between real and synthetic performance (i.e., reduce bias $\Delta=\theta^{s}-\theta^{r}$). Repeat from Step 2.
@@
    If REF remains unsatisfied, quantify fidelity limit by determining the smallest $\epsilon$ for which REF is certified. The new $(\epsilon,\alpha)$ reflects the best REF we can estimate with the available data. If new $\epsilon$ is large, this may indicate bias in our initial limited real-world sample, or persisting problems with the simulator.
@@
    Scale-up simulation testing with certified simulators.
    To use simulated results to assess the AV, we still need evidence that the simulator matches real operation, not just the limited sample we have. This requires prior evidence validating the simulator’s accuracy, as we explain later
  8. 8.
    Monitor REF after the AV deployment when more real-world data is collected. Recertify REF if it is violated.
We note that REF relies on a limited initial real-world sample. This means that a pfs estimate from this sample may differ from the true pfs. In particular, with very small pfs, such inaccuracy will be the norm, as a small sample will generally include zero failures. Within such a “biased” sample, REF may “certify” a simulator that happens to be affected by similar inaccuracy. Failing to pass a REF criterion is good indication that the simulator is inaccurate or the sample “biased”, and thus aids decisions at steps 4 and 5 above. To prove instead that the simulator is faithful to the real world, evidence of the appropriateness of physical models and quality of development need to be combined with the statistical observations (as we mentioned at step 7): a Bayesian argument that is the goal of future work. Step 8 with more real-world data will further verify whether the certified-REF remains valid for the actual AVs—not just for the initial limited sample.
@@
    Step 4: Doubling $t_{s}$ to 4000 with $k_{s}=102$ yields $\Pr(|\hat{\Delta}|\leq 0.02)\approx 0.91<0.95$; $(0.02,0.05)$-REF still uncertified.
@@
    $\hat{\Delta}=-0.005,\,\,\sigma_{\hat{\Delta}}\approx 0.0089,\,\,\Pr(|\hat{\Delta}|\leq 0.02)>0.95,$
    so $(0.02,0.05)$-REF is certified.
@@
How to incorporate prior knowledge about the simulator’s accuracy in Step 7, and how to integrate additional real-world evidence from Step 8 for estimating $\theta^{r}$, using both simulated and real-world data, requires proper Bayesian modelling. In practice, trust in a simulator will generally depend on quality of the physical models implemented, of the development, etc; leading to a Bayesian argument for confirming or weakening the initial trust through simulation results. More work is required also to address limitations of our example, e.g. that if failures are rare the sample means can no longer be approximated by normal distributions. Our intent here is just to demonstrate that formal probabilistic treatment of simulation accuracy is not just desirable, but likely feasible, rather than provide a comprehensive formal treatment.
A future direction is to define the mentioned Bayesian version of REF, where beliefs about the actual $\theta^{r}$ and $\theta^{s}$ are modeled, and being able to reason under rare failures. Once such a Bayesian REF is certified (e.g., with 95% probability that $\theta^{r}$ is no worse than $\theta^{s}$), existing models for “No Worse Than” reasoning [36, 40, 39, 43, 42] can be applied to support safety claims based on both real and simulated data.
@@ V Discussion and Conclusion
While we introduce pfs in our proof-of-concept models for scenario testing evidence, is pfs truly the best metric? Or should we consider higher abstraction level of “per trip” or “per mission”? What constitutes an acceptable pfs in AVs? Can we derive such thresholds from human-driven incident statistics at “per scenario” level? Answer to those question depends on how data are collected in practice for the specific AV under study, considering, e.g., if scenarios are i.i.d. [45] and if pfs is constant for the data collected.
@@
We caution that applying statistics in safety-critical domains requires great care: statistical inference, when used improperly, can be dangerously misleading. Thus, a solid statistical foundation is essential. Our goal is not to replace existing methods to AV safety, but to offer a complementary perspective: one that is driven by reasoning about the residual operational risk. By mapping challenges in AV testing to known problems in software reliability, we aim to bridge methodological gaps and enable more rigorous safety claims for AVs.